# Motion Control via Metric-Aligning Motion Matching

### v1

@@ Abstract.
Figure 1. Overview of our method. Our method aligns given original motion to given control sequence. Our method can take arbitrary control sequences, such as sketches, labels, audio, and motions. Our method solely relies on within-domain distance in original and control, without needing manual definition of mapping or training with annotated data.
[…]
@@ 1. Introduction
Early studies on motion-to-motion retargeting relied on hand-crafted mapping functions and strategies (Gleicher, 1998; Lee and Shin, 1999; Choi and Ko, 2000; Tak and Ko, 2005; Yamane et al., 2010; Rhodin et al., 2015; Seol et al., 2013). More recent approaches leverage large-scale learning-based frameworks to learn direct mappings between two motions (Villegas et al., 2018; Lim et al., 2019; Li et al., 2023b) or shared embeddings between two domains (Aberman et al., 2020; Lee et al., 2023; Li et al., 2024a). While deep learning techniques compute mappings empirically, they often require extensive annotated datasets and prolonged training to ensure robust generalization. Similarly, audio-driven motion synthesis (Alexanderson et al., 2023; Tseng et al., 2023; Li et al., 2024b) depends heavily on large paired datasets and substantial computational resources. Recent efforts to reduce data demands, as seen in works like GANimator (Li et al., 2022), GenMM (Li et al., 2023a), and SinDDM (Raab et al., 2023), demonstrate that motions can be generated from limited data. However, these methods still lack robust, built-in alignment strategies for complex temporal inputs and often require domain-specific constraints or supplementary annotations to measure cross-domain correspondences.
In this study, we propose a motion alignment technique that relies solely on within-domain distances, which can be naturally defined without requiring hand-crafted or learned mapping functions (Fig. 1). Our approach first computes distances among patches within the original motion and control sequences separately. We then search for couplings between patches in the control sequence and those in the original motion such that the within-domain distance structure is preserved as much as possible. These couplings are used to compute a weighted blend of original patches, resulting in an aligned motion. Technically, our method uses metric alignment techniques based on the fused semi-unbalanced Gromov-Wasserstein (FSUGW) optimization objective (Xu and Gould, 2024), a modified version of the Gromov-Wasserstein optimal transport (GW OT) problem. GW OT has been successfully applied in various applications, such as identifying correspondences between shapes (Solomon et al., 2016). By leveraging advancements in GW OT algorithms, our method achieves cross-domain coupling without explicitly computing cross-domain distances.
@@ 2.1. Motion Editing
Motion editing techniques aim to modify existing motion data or create new motion that satisfies new spatiotemporal constraints, adheres to specific stylistic requirements, or adapts to different environmental contexts. Early foundational approaches introduced the concept of spacetime constraints, formulating motion editing as an optimization problem guided by physical laws (Witkin and Kass, 1988; Gleicher, 1997). This line of research enabled animators to enforce physically plausible edits on motion sequences while maintaining user control over keyframed poses.
Subsequent works focused on lowering the barrier for nonexperts and improving workflow efficiency through sketch-based interfaces and other intuitive user controls (Thorne et al., 2004; Terra and Metoyer, 2004; Guay et al., 2015; Choi et al., 2016; Garcia et al., 2019). These methods allowed users to apply edits by sketching curves or gestures, which were then mapped to character poses and motion adjustments. Another approach involved creating poses by interpolating keyframes spatially (Igarashi et al., 2005). Parallel efforts explored physically guided animation tools, where motion was fine-tuned by automatically inferring or imposing physical properties, resulting in more believable outputs (Shapiro and Lee, 2011; Hahn et al., 2012).
More recent solutions integrate optimization frameworks directly into keyframe-based workflows, enabling artists to selectively refine motion segments without fully discarding manual control (Koyama and Goto, 2018; Ciccone et al., 2019). In addition, tangible or actuated puppets provide haptic feedback and a physical interface, further bridging digital and real-world motion editing (Jacobson et al., 2014; Yoshizaki et al., 2011). User performance itself has also served as an intuitive control mechanism (Dontcheva et al., 2003; Lockwood and Singh, 2012).
@@ 2.2. Motion Retargeting and Alignment
Motion retargeting transfers motion data from a source character to a target character, even when their skeletal structures or proportions differ. Early methods formulated this as an optimization problem, incorporating constraints such as low-level motion representations (Gleicher, 1998), kinematic constraints (Lee and Shin, 1999), end-effector trajectories (Choi and Ko, 2000), or dynamic feasibility (Tak and Ko, 2005). However, these approaches struggled with significant skeletal differences. Later approaches introduced partial correspondences (Yamane et al., 2010; Rhodin et al., 2015; Seol et al., 2013) or learned mappings independent of skeletal structures (Rhodin et al., 2014) to handle differing topologies.
The advent of deep learning and large motion capture datasets led to data-driven methods (Villegas et al., 2018; Lim et al., 2019; Aberman et al., 2020; Li et al., 2023b, 2024a). Early learning-based techniques leveraged cycle-consistency and adversarial losses (Villegas et al., 2018) or skeleton-aware components (Aberman et al., 2020). ACE (Li et al., 2023b) advanced retargeting for characters with significant skeletal differences using pretrained motion priors. Learning correspondences between general dynamical systems has also been investigated (Kim et al., 2020).
WalkTheDog (Li et al., 2024a) introduced motion alignment via a learned discretized phase manifold, enabling shared embeddings for characters with different topologies, such as humans and dogs. Their method uses frequency-scaled motion matching but requires extensive training and large datasets.
@@ 2.3. Audio-Driven Motion Synthesis
Audio-driven motion generation (e.g., dance, gesture synthesis) generates character motion from acoustic features, a long-standing topic in computer animation. For a comprehensive review, see (Zhu et al., 2024; Nyatsanga et al., 2023).
Early studies on dance synthesis primarily focused on synchronizing dance motions with musical beats, rhythms, and intensity using hand-crafted mappings (Shiratori et al., 2006; Shiratori and Ikeuchi, 2008; Ofli et al., 2008) or employing example-based approaches and statistical models (Ofli et al., 2012; Fan et al., 2012; Fukayama and Goto, 2015). Recent advances in deep learning have enabled the generation of more expressive dance motions by leveraging sophisticated motion generation models (Li et al., 2021; Lee et al., 2019; Siyao et al., 2022; Bhattacharya et al., 2024), including diffusion models (Alexanderson et al., 2023; Tseng et al., 2023; Li et al., 2024b). Other approaches combine learned choreo-musical embeddings with motion graphs (Au et al., 2022; Chen et al., 2021).
Parallel efforts in speech-driven gesture synthesis have similarly aimed to produce natural co-speech motion. Earlier approaches relied on extensive manual rules (Cassell et al., 2001; Kopp et al., 2003; Lee and Marsella, 2006; Lhommet et al., 2015), while more recent methods use learning-based methods (Ferstl and McDonnell, 2018; Hasegawa et al., 2018; Takeuchi et al., 2017; Alexanderson et al., 2020; Yazdian et al., 2021; Wu et al., 2021). However, many of these methods depend heavily on large paired audio-motion datasets (Alexanderson et al., 2023; Li et al., 2021; Chen et al., 2021; Lee et al., 2019; Siyao et al., 2022; Bhattacharya et al., 2024; Fan et al., 2012) and are computationally demanding (Tseng et al., 2023; Li et al., 2024b; Alexanderson et al., 2023), limiting their adaptability to new characters or resource-limited environments. Our approach bypasses these challenges by employing fused semi-unbalanced Gromov-Wasserstein optimal transport, which reduces reliance on large datasets and hand-crafted features. This enables more efficient and flexible audio-to-motion synthesis suitable for a wide range of contexts.
@@ 2.4. Motion Synthesis from Small Data
Building on recent advances in image generation from a single or a few examples (Shaham et al., 2019; Granot et al., 2022; Kulikov et al., 2023), several studies have explored synthesizing motion using only a limited number of example sequences. For instance, GANimator (Li et al., 2022) employs generative adversarial networks to generate diverse motion patterns from a single instance. GenMM (Li et al., 2023a), in contrast, bypasses the training phase entirely, enabling rapid motion synthesis through a bidirectional similarity-based matching framework. SinDDM (Raab et al., 2023) further improves the quality of generated motion using diffusion models.
While these approaches demonstrate significant flexibility in conditioned motion synthesis tasks – such as inbetweening, trajectory control (Li et al., 2022, 2023a), and style transfer within the same skeletal structure (Raab et al., 2023) – they often lack inherent mechanisms to handle more complex motion control by sequences. Consequently, achieving desirable outcomes in such scenarios frequently requires additional annotations or manual interventions, which can increase both labor and computational costs.
@@ 2.5. Optimal Transport
Optimal transport (OT) has been widely applied across various domains, including machine learning, computer vision, and neuroscience (Montesuma et al., 2024). For instance, Elnekave and Weiss (Elnekave and Weiss, 2022) introduced methods for generating natural images by matching patch distributions. Advancements in Gromov-Wasserstein (GW) distance have enabled mappings between disparate domains, such as shape correspondence and image-to-color distribution matching (Solomon et al., 2016). Recent developments in Unbalanced and Fused GW (Sejourne et al., 2021; Chizat et al., 2018) have further expanded OT applications to areas such as aligning individual brains (Thual et al., 2022) and unsupervised action segmentation (Xu and Gould, 2024). We incorporate the fused unbalanced GW OT problem solver (Xu and Gould, 2024) as a core component of our framework, offering a unified and effective solution to sequence-conditioned motion control.
@@ 3.1. Motion Representation
We represent a motion sequence as a series of time-indexed poses, as in a previous study (Li et al., 2023a), where each pose is described by root displacement $V$ and joint rotation $R$ in character space. At time $t$, the pose is given by $[V_{t},R_{t}]$, where $V\in\mathbb{R}^{T\times 3}$ represents the changes in the root joint position compared to consecutive frames, and $R\in\mathbb{R}^{T\times J\times 3\times 3}$ describes the rotation of each joint in the character’s root space in matrix representation. Here, $J$ denotes the number of joints.
@@ 3.2. Metric-Aligning Motion Matching
In the actual algorithm, we define the problem as the minimization of the fused semi-unbalanced Gromov-Wasserstein (FSUGW) loss $L_{FSUGW}$. Our formulation of $L_{FSUGW}$, as presented in Equations (1), (2), and (3), along with the projected mirror-descent algorithm employed as a component of our framework, follows the formulations in previous work (Xu and Gould, 2024). By minimizing this objective via MAMM, $X^{\prime}$ is optimized as described. Our contribution lies in integrating coarse-to-fine optimization (Sec. 3.2.5) and the patch extraction and blending process into the optimization loop to ensure temporal consistency (Sec. 3.2.4). Fig. 3 provides an overview of the framework.
@@ 3.2.1. Motion Patch Extraction
To enable fine-grained optimization, we follow previous patch-based generation methods (Li et al., 2023a; Elnekave and Weiss, 2022) by splitting the control sequence $Y$, original motion $X$, and aligned motion $X^{\prime}$ into overlapping temporal patches with a stride of one frame. These patches are labeled as $\tilde{Y}$, $\tilde{X}$, and $\tilde{X^{\prime}}$, respectively. Note that $\tilde{Y}$ and $\tilde{X}$ remain fixed, while $\tilde{X^{\prime}}$ is updated at every step of the optimization loop ($\tilde{X^{\prime}}\leftarrow$ ExtractPatches$(X^{\prime})$ ). Measuring the FSUGW loss at the patch level ensures the capture of local temporal structures and ensures continuity.
Hereafter, we denote the length of the patch sequences as $L_{X}=|\tilde{X}|$ and $L_{Y}=|\tilde{Y}|=|\tilde{X^{\prime}}|$.
@@ 3.2.4. Fused Semi-Unbalanced Gromov-Wasserstein (FSUGW) Block
We combine the Wasserstein loss and GW loss to define and optimize $L_{FSUGW}$ as an instance of the entropic fused semi-unbalanced Gromov-Wasserstein problem (Xu and Gould, 2024):
@@
Here, we define $a=\boldsymbol{1}_{L_{Y}}/{L_{Y}},\quad b=\boldsymbol{1}_{L_{X}}/{L_{X}}$, where $L_{Y}$ and $L_{X}$ are the number of data (frames) in $\tilde{Y}$ (or $\tilde{X^{\prime}}$) and $\tilde{X}$, respectively. The term $H(T)=-\sum_{i,j}T_{i,j}log(T_{i,j})$ represents the soft entropy constraint, which facilitates optimization via the projected mirror-descent algorithm (Solomon et al., 2016; Xu and Gould, 2024). The soft marginal constraint, defined as the $D_{KL}$ term, encourages bi-directional fidelity of the distribution of the aligned motion $X^{\prime}$ to that of $X$. While soft constraints are imposed on one side, we enforce hard constraints on the other side of the marginals, $T\boldsymbol{1}=a$, ensuring that $T$ satisfies the specified marginals. Intuitively, this constraint ensures that every part of the aligned motion $X^{\prime}$ corresponds to some part of the original motion $X$. The effects of the parameters $\alpha$ and $\lambda$ will be discussed in Sec. 5.1.
Optimizing $L_{FSUGW}$ w.r.t $T$ and $X^{\prime}$ is similar to the (fused unbalanced) Gromov-Wasserstein Barycenter problem (Peyré et al., 2016; Thual et al., 2022), which is typically solved using a block coordinate descent algorithm. Inspired by this, our proposed FSUGW block optimizes $L_{FSUGW}$ by alternating the following steps:
@@
    Optimize $T$ with fixed $X^{\prime}$: Given $X^{\prime}$, we find $T$ as a local minimum of $L_{FSUGW}$ by applying the projected mirror-descent algorithm, identical to the version in (Xu and Gould, 2024) except for the parameters. At each stage, $T$ is initialized as $\mathbf{a}\mathbf{b}^{T}$, following previous practices.

### v2

@@ Abstract.
Figure 1. Overview of our method. Our method aligns given original motion to given control sequence. Our method can take arbitrary control sequences, such as sketches, labels, audio, and motions. Our method solely relies on within-domain distance in original and control, without needing manual definition of mapping or training with annotated data.: :
@@ 1. Introduction
Early studies on motion-to-motion retargeting relied on hand-crafted mapping functions and strategies (gleicher_retargetting_1998; lee_hierarchical_1999; choi_online_2000; tak_physically-based_2005; yamane_animating_2010; rhodin_generalizing_2015; seol_creature_2013). More recent approaches leverage large-scale learning-based frameworks to learn direct mappings between two motions (villegas_neural_2018; lim_pmnet_2019; li_ace_2023) or shared embeddings between two domains (aberman_skeleton-aware_2020; lee_same_2023; li_walkthedog_2024). While deep learning techniques compute mappings empirically, they often require extensive annotated datasets and prolonged training to ensure robust generalization. Similarly, audio-driven motion synthesis (alexanderson_listen_2023; tseng_edge_2023; li_lodge_2024) depends heavily on large paired datasets and substantial computational resources. Recent efforts to reduce data demands, as seen in works like GANimator (li_ganimator_2022), GenMM (li_example-based_2023), and SinDDM (raab_single_2023), demonstrate that motions can be generated from limited data. However, these methods still lack robust, built-in alignment strategies for complex temporal inputs and often require domain-specific constraints or supplementary annotations to measure cross-domain correspondences.
In this study, we propose a motion alignment technique that relies solely on within-domain distances, which can be naturally defined without requiring hand-crafted or learned mapping functions (Fig. 1). Our approach first computes distances among patches within the original motion and control sequences separately. We then search for couplings between patches in the control sequence and those in the original motion such that the within-domain distance structure is preserved as much as possible. These couplings are used to compute a weighted blend of original patches, resulting in an aligned motion. Technically, our method uses metric alignment techniques based on the fused semi-unbalanced Gromov-Wasserstein (FSUGW) optimization objective (xu_temporally_2024), a modified version of the Gromov-Wasserstein optimal transport (GW OT) problem. GW OT has been successfully applied in various applications, such as identifying correspondences between shapes (solomon_entropic_2016). By leveraging advancements in GW OT algorithms, our method achieves cross-domain coupling without explicitly computing cross-domain distances.
@@ 2.1. Motion Editing
Motion editing techniques aim to modify existing motion data or create new motion that satisfies new spatiotemporal constraints, adheres to specific stylistic requirements, or adapts to different environmental contexts. Early foundational approaches introduced the concept of spacetime constraints, formulating motion editing as an optimization problem guided by physical laws (witkin_spacetime_1988; gleicher_motion_1997). This line of research enabled animators to enforce physically plausible edits on motion sequences while maintaining user control over keyframed poses.
Subsequent works focused on lowering the barrier for nonexperts and improving workflow efficiency through sketch-based interfaces and other intuitive user controls (thorne_motion_2004; terra_performance_2004; guay_space-time_2015; choi_sketchimo_2016; garcia_spatial_2019). These methods allowed users to apply edits by sketching curves or gestures, which were then mapped to character poses and motion adjustments. Another approach involved creating poses by interpolating keyframes spatially (igarashi_spatial_2005). Parallel efforts explored physically guided animation tools, where motion was fine-tuned by automatically inferring or imposing physical properties, resulting in more believable outputs (shapiro_practical_2011; hahn_rig-space_2012).
More recent solutions integrate optimization frameworks directly into keyframe-based workflows, enabling artists to selectively refine motion segments without fully discarding manual control (koyama_optimo_2018; ciccone_tangent-space_2019). In addition, tangible or actuated puppets provide haptic feedback and a physical interface, further bridging digital and real-world motion editing (jacobson_tangible_2014; yoshizaki_actuated_2011). User performance itself has also served as an intuitive control mechanism (dontcheva_layered_2003; lockwood_finger_2012).
@@ 2.2. Motion Retargeting and Alignment
Motion retargeting transfers motion data from a source character to a target character, even when their skeletal structures or proportions differ. Early methods formulated this as an optimization problem, incorporating constraints such as low-level motion representations (gleicher_retargetting_1998), kinematic constraints (lee_hierarchical_1999), end-effector trajectories (choi_online_2000), or dynamic feasibility (tak_physically-based_2005). However, these approaches struggled with significant skeletal differences. Later approaches introduced partial correspondences (yamane_animating_2010; rhodin_generalizing_2015; seol_creature_2013) or learned mappings independent of skeletal structures (rhodin_interactive_2014) to handle differing topologies.
The advent of deep learning and large motion capture datasets led to data-driven methods (villegas_neural_2018; lim_pmnet_2019; aberman_skeleton-aware_2020; li_ace_2023; li_walkthedog_2024). Early learning-based techniques leveraged cycle-consistency and adversarial losses (villegas_neural_2018) or skeleton-aware components (aberman_skeleton-aware_2020). ACE (li_ace_2023) advanced retargeting for characters with significant skeletal differences using pretrained motion priors. Learning correspondences between general dynamical systems has also been investigated (kim_learning_2020).
WalkTheDog (li_walkthedog_2024) introduced motion alignment via a learned discretized phase manifold, enabling shared embeddings for characters with different topologies, such as humans and dogs. Their method uses frequency-scaled motion matching but requires extensive training and large datasets.
@@ 2.3. Audio-Driven Motion Synthesis
Audio-driven motion generation (e.g., dance, gesture synthesis) generates character motion from acoustic features, a long-standing topic in computer animation. For a comprehensive review, see (zhu_human_2024; nyatsanga_comprehensive_2023).
Early studies on dance synthesis primarily focused on synchronizing dance motions with musical beats, rhythms, and intensity using hand-crafted mappings (shiratori_dancing--music_2006; shiratori_synthesis_2008; ofli_audio-driven_2008) or employing example-based approaches and statistical models (ofli_learn2dance_2012; fan_example-based_2012; fukayama_2015_music). Recent advances in deep learning have enabled the generation of more expressive dance motions by leveraging sophisticated motion generation models (li_ai_2021; lee_dancing_2019; siyao_bailando_2022; bhattacharya_danceanyway_2024), including diffusion models (alexanderson_listen_2023; tseng_edge_2023; li_lodge_2024). Other approaches combine learned choreo-musical embeddings with motion graphs (au_choreograph_2022; chen_choreomaster_2021).
Parallel efforts in speech-driven gesture synthesis have similarly aimed to produce natural co-speech motion. Earlier approaches relied on extensive manual rules (cassell_beat_2001; kopp_max_2003; lee_nonverbal_2006; lhommet_cerebella_2015), while more recent methods use learning-based methods (ferstl_investigating_2018; hasegawa_evaluation_2018; takeuchi_speech--gesture_2017; alexanderson_style-controllable_2020; yazdian_gesture2vec_2021; wu_probabilistic_2021). However, many of these methods depend heavily on large paired audio-motion datasets (alexanderson_listen_2023; li_ai_2021; chen_choreomaster_2021; lee_dancing_2019; siyao_bailando_2022; bhattacharya_danceanyway_2024; fan_example-based_2012) and are computationally demanding (tseng_edge_2023; li_lodge_2024; alexanderson_listen_2023), limiting their adaptability to new characters or resource-limited environments. Our approach bypasses these challenges by employing fused semi-unbalanced Gromov-Wasserstein optimal transport, which reduces reliance on large datasets and hand-crafted features. This enables more efficient and flexible audio-to-motion synthesis suitable for a wide range of contexts.
@@ 2.4. Motion Synthesis from Small Data
Building on recent advances in image generation from a single or a few examples (shaham_singan_2019; granot_drop_2022; kulikov_sinddm_2023), several studies have explored synthesizing motion using only a limited number of example sequences. For instance, GANimator (li_ganimator_2022) employs generative adversarial networks to generate diverse motion patterns from a single instance. GenMM (li_example-based_2023), in contrast, bypasses the training phase entirely, enabling rapid motion synthesis through a bidirectional similarity-based matching framework. SinDDM (raab_single_2023) further improves the quality of generated motion using diffusion models.
While these approaches demonstrate significant flexibility in conditioned motion synthesis tasks – such as inbetweening, trajectory control (li_ganimator_2022; li_example-based_2023), and style transfer within the same skeletal structure (raab_single_2023) – they often lack inherent mechanisms to handle more complex motion control by sequences. Consequently, achieving desirable outcomes in such scenarios frequently requires additional annotations or manual interventions, which can increase both labor and computational costs.
@@ 2.5. Optimal Transport
Optimal transport (OT) has been widely applied across various domains, including machine learning, computer vision, and neuroscience (Montesuma_recent_2024). For instance, Elnekave and Weiss (elnekave_generating_2022) introduced methods for generating natural images by matching patch distributions. Advancements in Gromov-Wasserstein (GW) distance have enabled mappings between disparate domains, such as shape correspondence and image-to-color distribution matching (solomon_entropic_2016). Recent developments in Unbalanced and Fused GW (sejourne_unbalanced_2021; chizat_scaling_2018) have further expanded OT applications to areas such as aligning individual brains (thual_aligning_2022) and unsupervised action segmentation (xu_temporally_2024). We incorporate the fused unbalanced GW OT problem solver (xu_temporally_2024) as a core component of our framework, offering a unified and effective solution to sequence-conditioned motion control.
@@ 3.1. Motion Representation
We represent a motion sequence as a series of time-indexed poses, as in a previous study (li_example-based_2023), where each pose is described by root displacement $V$ and joint rotation $R$ in character space. At time $t$, the pose is given by $V_{t},R_{t}]$, where $V\in\mathbb{R}^{T\times 3}$ represents the changes in the root joint position compared to consecutive frames, and $R\in\mathbb{R}^{T\times J\times 3\times 3}$ describes the rotation of each joint in the character’s root space in matrix representation. Here, $J$ denotes the number of joints.
@@ 3.2. Metric-Aligning Motion Matching
In the actual algorithm, we define the problem as the minimization of the fused semi-unbalanced Gromov-Wasserstein (FSUGW) loss $L_{FSUGW}$. Our formulation of $L_{FSUGW}$, as presented in Equations (1), (2), and (3), along with the projected mirror-descent algorithm employed as a component of our framework, follows the formulations in previous work (xu_temporally_2024). By minimizing this objective via MAMM, $X^{\prime}$ is optimized as described. Our contribution lies in integrating coarse-to-fine optimization (Sec. 3.2.5) and the patch extraction and blending process into the optimization loop to ensure temporal consistency (Sec. 3.2.4). Fig. 3 provides an overview of the framework.
@@ 3.2.1. Motion Patch Extraction
[…]
@@ 3.2.4. Fused Semi-Unbalanced Gromov-Wasserstein (FSUGW) Block
We combine the Wasserstein loss and GW loss to define and optimize $L_{FSUGW}$ as an instance of the entropic fused semi-unbalanced Gromov-Wasserstein problem (xu_temporally_2024):
@@
Here, we define $a=\boldsymbol{1}_{L_{Y}}/{L_{Y}},\quad b=\boldsymbol{1}_{L_{X}}/{L_{X}}$, where $L_{Y}$ and $L_{X}$ are the number of data (frames) in $\tilde{Y}$ (or $\tilde{X^{\prime}}$) and $\tilde{X}$, respectively. The term $H(T)=-\sum_{i,j}T_{i,j}log(T_{i,j})$ represents the soft entropy constraint, which facilitates optimization via the projected mirror-descent algorithm (solomon_entropic_2016; xu_temporally_2024). The soft marginal constraint, defined as the $D_{KL}$ term, encourages bi-directional fidelity of the distribution of the aligned motion $X^{\prime}$ to that of $X$. While soft constraints are imposed on one side, we enforce hard constraints on the other side of the marginals, $T\boldsymbol{1}=a$, ensuring that $T$ satisfies the specified marginals. Intuitively, this constraint ensures that every part of the aligned motion $X^{\prime}$ corresponds to some part of the original motion $X$. The effects of the parameters $\alpha$ and $\lambda$ will be discussed in Sec. 5.1.
Optimizing $L_{FSUGW}$ w.r.t $T$ and $X^{\prime}$ is similar to the (fused unbalanced) Gromov-Wasserstein Barycenter problem (peyre_gromov-wasserstein_2016; thual_aligning_2022), which is typically solved using a block coordinate descent algorithm. Inspired by this, our proposed FSUGW block optimizes $L_{FSUGW}$ by alternating the following steps:
@@
    Optimize $T$ with fixed $X^{\prime}$: Given $X^{\prime}$, we find $T$ as a local minimum of $L_{FSUGW}$ by applying the projected mirror-descent algorithm, identical to the version in (xu_temporally_2024) except for the parameters. At each stage, $T$ is initialized as $\mathbf{a}\mathbf{b}^{T}$, following previous practices.